OpenStudy (anonymous):

Cassy shoots a large marble (Marble A, mass: 0.06 kg) at a smaller marble (Marble B, mass: 0.03 kg) that is sitting still. Marble A was initially moving at a velocity of 0.7 m/s, but after the collision it has a velocity of –0.2 m/s. What is the resulting velocity of marble B after the collision? Be sure to show your work for solving this problem along with the final answer.

OpenStudy (anonymous):

total momentum of system should be conservative so Ma*Va=total momentum so 0.06*0.7=0.0042 after sollison 0.06*-0.2=-0.0012 so 0.0042-0.0012=0.003 is momentum of marbel B so its velocity= 0.003/0.03

OpenStudy (anonymous):

so i would divide 0.003/0.03 for my answer

OpenStudy (anonymous):

no it would be 1m/s

OpenStudy (anonymous):

yea u divided to get that right ?